OpenStudy (thedj4jc):

Simplify The square root of 125 A. 5 times the square root of 5 B. the square root of 5 C. the square root of 5 divided by 5/over 5

OpenStudy (thedj4jc):

@ʕ•́ᴥ•̀ʔっ

OpenStudy (freethinker):

break down 125

OpenStudy (anonymous):

5 is the square root of 5

OpenStudy (freethinker):

25 * 5

OpenStudy (freethinker):

\(\large \sqrt{25 \times 5}\) SIMPLIFY

OpenStudy (thedj4jc):

That'd be 5 * the square root of 5! Thanks you!!!

OpenStudy (anonymous):

Yeah, we find the perfect square factors.

OpenStudy (anonymous):

One of them is 25. 125 / 25 = 5 So we have: \(\sf \sqrt{25} \times \sqrt{5}\) Simplify the square root of 25: \(\sf 5 \times \sqrt{5}\)

OpenStudy (anonymous):

Bravo!